What is the symbol for probability of success in a binomial trial?
The symbol for probability of success in a binomial trial is the letter p. It is the symbol used for probability in all statistical testing.
Two coins are tossed. What is the probability of tossing at least one head?
There are 2 coins, 1 and 2. Each has two possibilities, H or T.
The possibilities are:
1H, 2H
1H, 2T
1T, 2H
1T, 2T
Each possibility has an equal chance of happening.
The chance of tossing at least one head is 3/4.
If two events are mutually exclusive then they must be dependent?
No, if two events are mutually exclusive, they cannot both occur. If one occurs, it means the second can not occur.
What is the difference between a normal distribution and the standard normal distribution?
The standard normal distribution is a normal distribution with mean 0 and variance 1.
If you sit on a moving bus and toss a coin straight up into the air where will it land?
The coin is traveling at the same velocity as the tosser, so it will land the same as if you were standing still.
A six-sided cube is rolled Find the probability of rolling a six?
Probability = (number of successful outcomes) / (number of possible outcomes)
Possible outcomes: 6
Successful outcomes: 1
Probability = 1/6 = 16 and 2/3 percent.
odds
"The odds against an event is a ratio of the probability that the event will fail to occur (failure) to the probability that the event will occur (success). To find odds you must first know or determine the probability of success and the probability of failure.
Odds against event = P(event fails to occur)/P(event occurs) = P(failure)/P(success)
The odds in favor of an event are expressed as a ratio of the probability that the event will occur to the probability that the event will fail to occur.
Odds in favor of event = P(event occurs)/P(event fails to occur) = P(success)/P(failure)"
Allen R. Angel, Christine D. Abbott, Dennis C. Runde. A Survey of Mathematics with Applications. Pearson Custom Publishing 2009. Pages 286-288.
What is the probability that the number 23 will be randomly picked out of 100?
Assuming then that there are 100 numbers, 1-100, the probability of the number 23 being randomly picked out of 100 is: 1/100 or 0.01.
What are the odds for drawing a King from a regular deck of 52 cards?
There are 4 kings and 52 cards so the probability of drawing a king are 4/52 or 2/26.
What is the area under the normal distribution curve to the right of z that equals 3.24?
0.0006 (approx).
Probability of not blue is the probability of white. The probability of white is 11/(11+21) or 11/32.
The probability that 2 people have the same number is 2 out of 10
A multiple choice question has has 5 choices What is the probability of guessing it correctly?
There is 1 right answer out of 5 possible answers, so the probability of guessing it correctly is 1/5 or 20% or 0.2.
A spinner a-f what is the probability of the spinner landing on b then landing on a?
Assuming that it is a regular shaped spinner, the probability is 1/6*1/6 = 1/36
If you roll a die what is the probability that a three will appear on top?
I think it's 1 out of 6 chance.
Why is the probability measured as a number between zero and one?
Probability is the ratio of no. of possible outcomes to the maximum no. of outcomes .
If any event doesnt occurs ,then it is 0/N,where N=maximum no of outcomes .and that is 0 itself ,hence that is signified by zero ,or will at max, all the possible outcomes wil occur ,which means N/N that is denoted by 1.
Hence probabilty for the occurance for any event is between 0 and 1.
If you toss a number cube 600 times what is the most probable number of times you will get a 3?
100/600 or 1/6
If 20 newborn babies are randomly selected how many different gender sequences are possible?
You have 2 choices for 20 times, so the number of sequences is 220 or 1048576.
Relative frequency approximation is conducting experiments and counting the number of times the event occurs divided by the total number of events. The classical approach is determine the number of ways the event can occur divided by the total number of events.
What did the psychiastrist say to the guy who thought he was a deck of cards?
I will deal with you later
Why does a researcher want to go from a normal distribution to a standard normal distribution?
A researcher wants to go from a normal distribution to a standard normal distribution because the latter allows him/her to make the correspondence between the area and the probability. Though events in the real world rarely follow a standard normal distribution, z-scores are convenient calculations of area that can be used with any/all normal distributions. Meaning: once a researcher has translated raw data into a standard normal distribution (z-score), he/she can then find its associated probability.
What is the probability of drawing a spade from a deck of cards?
A standard deck is 52 cards. These 52 cards are divided evenly into 4 suits - spades, clubs, diamonds, and hearts. There are 13 cards in each suit, ace through king. Therefore, there are 13 spades in a deck of 52 cards. 13 is one fourth of 52 which makes the probability of drawing a spade 1/4, .25, or 25 percent.
How many different four-letter initials are there that begin with A?
26^3. There are 26 possible letters, so that is why we multipy by 26. Because there are 4 positions, and the first is taken with "A", that only leaves 3 positions left. As there are 26 possible letters in each position, you multiply by 26 three times, or, 26 to the third power.